OpenStudy (anonymous):

using f(p)= 180-0.3p^2 where p is the price in dollars and f(p) is the number of items sold. at what price will maximum revenue be generated? rounded to the nearest cent

OpenStudy (anonymous):

that's weird..i got a maximum turning point at p=0 and f(p)=180

OpenStudy (anonymous):

I think i did it wrong, ill check

OpenStudy (anonymous):

okay thank you

OpenStudy (anonymous):

I got p=0 again, what's the answer though?

OpenStudy (anonymous):

$0 just looks wrong...so i dont know

OpenStudy (anonymous):

If something is free 180 items will be sold. Lol.

OpenStudy (anonymous):

I just did f'(p) = -0.6p When f'(p)=0, p=0 and f(p)=180 And then i found the nature of the turning point... and it's a max but you're trying to generate max revenue? why is everything gonna be free then

OpenStudy (anonymous):

f(p) should be revenue. As we are trying to maximize this function. p should be items sold. I think....

OpenStudy (anonymous):

yeah that's what puzzled me...ballards are you sure you typed out the Q correctly

OpenStudy (anonymous):

here it is again straight from the test im trying to correct

OpenStudy (anonymous):

The demand function for a certain comodity is given by f(p)= 180 - 0.3p^2, where p is the price on dollars and f(p) is the number of items sold. #d) at what price will maximum revenue be generated? Round to the nearest cent

OpenStudy (anonymous):

I think p is the price you pay to manufacture the item... so if you pay $0 then you still manage to sell 180 items and thus you generate max. revenue? Considering that you've made no loss? I don't know, ha.

OpenStudy (anonymous):

ha ha ha thanks for trying though

OpenStudy (anonymous):

Do you have answers for it the sheet you're doing?

OpenStudy (anonymous):

for the*

OpenStudy (anonymous):

no it was a test we have to correct for partial credit

OpenStudy (anonymous):

this is what i did and she only counted off 3 instead of 5

OpenStudy (anonymous):

162 = 180-.3p^2 -180 -180 -18= -.3p^2 /-.3 /-.3 60=p^2 p=7.75